OpenStudy (anonymous):

How many different arrangements of the letters in the word PARALLEL are there?

OpenStudy (anonymous):

if all the letters were different it would be \(8!\)

OpenStudy (anonymous):

56 3,360 40,320

OpenStudy (anonymous):

but since you have 2 "A" that you cannot tell apart, and also three identical "L" it is \[\frac{8!}{2!3!}\]

OpenStudy (anonymous):

do you know how to compute this number?

OpenStudy (anonymous):

Im confused

OpenStudy (anonymous):

ok lets go slow do you know what \(8!\) means ?

OpenStudy (anonymous):

ya 8*7*6*5*4*3*2*1

OpenStudy (anonymous):

if you have 8 different letters, or symbols or whatever, the number of ways they can be arranges is \(8!\) so you got that part yes?

OpenStudy (anonymous):

ok good, but in this case you have some letters that you cannot tell apart

OpenStudy (anonymous):

Okay im not sure how to do this srry

OpenStudy (anonymous):

there are two "E" and the number of ways to order the 2 "E" is 2

OpenStudy (anonymous):

and also there are 3 "L" and the number of ways to order the 3 "L" are \(3!=3\times 2=6\)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so to answer your question, you have to divide \[\frac{8!}{2!3!}=\frac{8\times 7\times 6\times 5\times 4\times 3\times 2}{2\times 3\times 2}\]

OpenStudy (anonymous):

cancel first, multiply last

OpenStudy (anonymous):

you can compute \[8\times 7\times 6\times 5\times 2\] for example

OpenStudy (anonymous):

Okay so it will be 6720

OpenStudy (anonymous):

still probably need a calculator i get \(3360\) http://www.wolframalpha.com/input/?i=8*7*6*5*2

OpenStudy (anonymous):

or just type it in directly and don't bother with the cancellation still 3360 http://www.wolframalpha.com/input/?i=8!%2F%282!3!%29

OpenStudy (anonymous):

O okay